OpenStudy (anonymous):

i need help how do i find C in 2/3(c-18)=7

OpenStudy (anonymous):

2/3(c-18)=17 (3/2)(2/3(c - 18)) = (3/2)17 c - 18 = 51/2 c = 51/2 + 36/2 = 87/2

OpenStudy (anonymous):

that is how you can fine c

OpenStudy (anonymous):

c =43.5 ??

OpenStudy (anonymous):

yea it does

OpenStudy (anonymous):

heey what algebraic expression for the phrase the sum of g and 3.

OpenStudy (anonymous):

did u get the answer already

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ook